solve the system of equations by substitution
5x+y=-4
y=-x

Found 2 solutions by rfer, mananth:
Answer by rfer(16322)

5x-x=-4
4x=-4
x=-1
------
5*-1+y=-4
-5+y=-4
y=+1
-------------
proof
y=-x
1=--1

Answer by mananth(16949)

5 x + 1 y = -4 .............1

1 y = 0 - -1 x
Plug the value of y in (1)
5x+1(0- x)=-4
5x-x)=-4
5x-1x=-4 0
4x=-4
/4
x=-1
Plug the value of x in (1)
5x+y=-4
5*-1+y=-4
-5+y=-4
1y=1
/1
y=1